# Symbol letters of Feynman integrals from Gram determinants

### v2

@@ II Setup and rational letters
According to the findings in Jiang and Yang (2023), an integral family is not associated with just a single BR, but rather a tree of BRs that form a recursive structure. Each BR is defined by a specific $u$-function in Eq. (3). The so-called standard BR, which includes all Baikov variables in $u$, serves as the topmost block in the recursive structure. It is the super-BR of all other BRs. Given a BR, one can derive its sub-BRs by sequentially integrating out variables. In the case of the $\{1,2,3\}$ top-sector and its sub-sectors in the sunrise family, the recursive structure is illustrated in Fig. 2. The arrows in the diagram are labeled with the variables to be integrated out. The blue blocks represent zero-BRs, where all integrals are scaleless and vanish. Each sector can possess several BRs. For instance, the $\{1,2,3\}$ sector can be expressed in terms of the $u_{1}$, $u_{2}$ and the standard BRs. Among them, $u_{1}$ and $u_{2}$ are the lowest ones in their branches, i.e., they don’t have descendants that can be used to represent the sector $\{1,2,3\}$. We will refer to them as the “minimal BRs” for this sector. This concept will play an important role in our construction. It is typical that a genuine multi-loop sector has more than one minimal BRs. A “contra-example” is the $\{1,3\}$ sector, which can be represented by the $u_{3}$-BR and all its super-BRs as depicted in Fig. 2. It can be seen that $u_{3}$ is the only minimal BR for this sector. This reflects the fact that integrals in the $\{1,3\}$ sector are products of one-loop integrals.
@@ III Bootstrapping algebraic letters
where $\bm{k}$ represents a sequence of momenta, while $q_{i}$ and $q_{j}$ denote two momenta not present in $\bm{k}$. For convenience, we allow $\bm{k}$ to be an empty sequence, and define $G()=1$. These Grams are linked by the identity $B^{2}+DE=AC$ Dlapa et al. (2021); Chen et al. (2022b), which leads to
@@
We categorize all the Grams in the family based on their powers $\alpha+\beta\epsilon$ in the $u$-functions. We have $\mathcal{G}_{1}$ containing those with $\alpha\in\mathbb{Z}$ and $\mathcal{G}_{2}$ containing those with $\alpha+1/2\in\mathbb{Z}$. We require that either $A,C\in\mathcal{G}_{2}$ or $D,E\in\mathcal{G}_{1}$. Furthermore, we require that $\sqrt{\hat{A}\hat{C}}$ is a product of the algebraic LS obtained in the previous section. We also demand that $\hat{A}\hat{C}$ is not a perfect square, otherwise it would result in a rational letter.
@@ V Summary and outlook
Our method can be applied to non-planar topologies as well. For non-planar topologies, a complication is that there are usually more spurious letters appearing in our predictions. One way to reduce the number of spurious letters is to choose the ISPs appropriately, such that all propagator variables appear at most quadratically in the $u$-function of the standard representation. This can often be achieved by choosing the ISPs to be linear in loop momenta. Such choices simplifies the recursive structure of BRs and also simplifies the calculations. We elaborate more on these details in an upcoming article Jiang et al. .

### v3

@@ II Setup and rational letters
According to the findings in Jiang and Yang (2023), an integral family is not associated with just a single BR, but rather a tree of BRs that form a recursive structure. Each BR is defined by a specific $u$-function in Eq. (3). The so-called standard BR, which includes all Baikov variables in $u$, serves as the topmost block in the recursive structure. It takes the form:
$I_{\text{std}}=\frac{\pi^{(L-N)/2}\det(A^{a}_{ij})}{\prod_{i=1}^{L}\Gamma\left(\frac{d-K+i}{2}\right)}\left[G(p_{1},\ldots,p_{E})\right]^{-(d-E-1)/2}\\ \times\int\frac{\mathrm{d}x_{1}\cdots\mathrm{d}x_{N}}{x_{1}^{a_{1}}\cdots x_{N}^{a_{N}}}\left[P^{L}_{N}(x_{1},\cdots,x_{N})\right]^{(d-K-1)/2}\,,$ (4)
where $E$ is the number of independent external legs, $L$ is the number of loops, and $K\equiv L+E$. The Baikov polynomial $P^{L}_{N}$ is given by
$P^{L}_{N}(x_{1},\ldots,x_{N})=G(q_{1},q_{2},\ldots,q_{K})\,,$ (5)
where $\{q_{1},\ldots,q_{K}\}$ is the set of all loop and independent external momenta. The scalar products $q_{i}\cdot q_{j}$ can then be expressed as linear combinations of the Baikov variables. The standard BR is the super-BR of all other BRs. Given a BR, one can derive its sub-BRs by sequentially integrating out variables. It is proven in Jiang and Yang (2023) that such variables can always be found. They are quadratic in the polynomial and can be integrated out using the general formula
$\int_{c_{1}}^{c_{2}}z^{n}\left[-A(z-c_{1})(z-c_{2})\right]^{\gamma}\mathrm{d}z=(-A)^{\gamma}\,(c_{2}-c_{1})^{1+2\gamma}$ (6)
$\frac{\Gamma(1\!+\!\gamma)^{2}}{\Gamma(2(1\!+\!\gamma))}\!\!\left(\!\!\frac{c_{1}+c_{2}}{2}\!\!\right)^{n}\!\!\!{}_{2}F_{1}\left(-\frac{n}{2},\!\frac{1\!-\!n}{2},\!\frac{3}{2}\!+\!\gamma;\left(\!\frac{c_{1}\!-\!c_{2}}{c_{1}\!+\!c_{2}}\!\right)^{2}\!\right).$
When $n\geq 0$, the hypergeometric function ${}_{2}F_{1}$ truncates to a polynomial. In particular, when $n=0,1$, it equals to 1. This integration can be performed recursively due to the structure of Gram determinants, thus all the BRs form a tree-and-net structure. In the case of the $\{1,2,3\}$ top-sector and its sub-sectors in the sunrise family, the recursive structure is illustrated in Fig. 2. The arrows in the diagram are labeled with the variables to be integrated out. The blue blocks represent zero-BRs, where all integrals are scaleless and vanish. Each sector can possess several BRs. For instance, the $\{1,2,3\}$ sector can be expressed in terms of the $u_{1}$, $u_{2}$ and the standard BRs. Among them, $u_{1}$ and $u_{2}$ are the lowest ones in their branches, i.e., they don’t have descendants that can be used to represent the sector $\{1,2,3\}$. We will refer to them as the “minimal BRs” for this sector. This concept will play an important role in our construction. It is typical that a genuine multi-loop sector has more than one minimal BRs. A “contra-example” is the $\{1,3\}$ sector, which can be represented by the $u_{3}$-BR and all its super-BRs as depicted in Fig. 2. It can be seen that $u_{3}$ is the only minimal BR for this sector. This reflects the fact that integrals in the $\{1,3\}$ sector are products of one-loop integrals.
@@ III Bootstrapping algebraic letters
where $\bm{k}$ represents a sequence of momenta, while $q_{i}$ and $q_{j}$ denote two momenta not present in $\bm{k}$. Note that $\bm{k}$ $q_{i}$ and $q_{j}$ can be loop momenta. The five Grams are not chosen arbitrarily, but are taken from the recursive structure of the BRs. Their forms are determined once we specify how the variables are integrated out from the topmost BR. For convenience, we also allow $\bm{k}$ to be an empty sequence, and define $G()=1$. These Grams are linked by the identity $B^{2}+DE=AC$ Dlapa et al. (2021); Chen et al. (2022b), which leads to
@@
We categorize all the Grams in the family based on their powers $\alpha+\beta\epsilon$ in the $u$-functions. We define the sets $\mathcal{G}_{1}$ containing those with $\alpha\in\mathbb{Z}$ and $\mathcal{G}_{2}$ containing those with $\alpha+1/2\in\mathbb{Z}$. We require that either $A,C\in\mathcal{G}_{2}$ or $D,E\in\mathcal{G}_{1}$. Furthermore, we require that $\sqrt{\hat{A}\hat{C}}$ is a product of the algebraic LS obtained in the previous section. We also demand that $\hat{A}\hat{C}$ is not a perfect square, otherwise it would result in a rational letter.
@@ V Summary and outlook
Our method can be applied to non-planar topologies as well. For non-planar topologies, a complication is that there are usually more spurious letters appearing in our predictions. One way to reduce the number of spurious letters is to choose the ISPs appropriately, such that all propagator variables appear at most quadratically in the $u$-function of the standard representation. This can often be achieved by choosing the ISPs to be linear in loop momenta. Such choices simplify the recursive structure of BRs and also simplify the calculations. We elaborate more on these details in an upcoming article.